Explain how the length contraction and time dilation formulas might be used to indicate that c is the limiting speed in the universe.



1
Expert's answer
2021-05-13T09:20:20-0400

The equation for length contraction:


"L=L_0\sqrt{1-(v/c)^2}"

The equation for time dilation:


"t=\frac{t_0}{\sqrt{1-(v/c)^2}}."

If something moves at the speed of light, the length becomes 0 and the time becomes infinity. So, no signal or interaction can move faster than the speed of light because they will simply not be able to occur.